DOTSEQN: Flush-left equations with dotted leaders to the equation number

Donald Arseneau

version v1.1, revised 1995/03/22

1 Introduction

The dotseqn package provides a different format for typesetting equations, one reportedly used in ‘old style Britsh books’ – equations aligned on the left, with dots on the right leading to the equation number. Like this

x = \frac{−b ± \sqrt{b^{2} − 4ac}}{2a} . . . . . . . . . . . . . . . . . . . . (1)

Equations without numbers have no leaders.

This package behaves a lot like fleqn (so that option is superfluous). In particular, the equations are not centered, but indented a constant distance from the left margin, controlled by \mathindent. The default indentation is the same as for lists.

2 Affected environments

Math environments with equation numbers, equation and eqnarray, are changed to produce left-justified equations, and to draw dotted leaders between the equation and the equation number. If there is no number, as specified by \nonumber, then no leaders are drawn.

Other math environments – displaymath, eqnarray∗, and \[ … \] – do not produce equation numbers and therefore behave the same as they do under the fleqn document option.

3 Options

This package is very similar to the fleqn document class option, so specifying fleqn is superfluous. The leqno class option is forbidden because it is incompatible with the dotted style of equation labeling. (Where would the leaders go?) As is the case for the ordinary fleqn option, the equation indentation is controlled by the length \mathindent, with a default indentation equal to the indentation of lists. It can be changed in the document preamble (with \setlength). Reasonable alternatives are zero (0pt) or the paragraph indentation (\parindent).

The dotseqn package has two options of its own which can be specified with the \usepackage command. The leftjust option causes the first column in eqnarray and eqnarray∗ to be flush left. Ordinarily, these pieces are flush right against the middle column.

Another option is nocolsep, which gets rid of that (ridiculous) extra space in the second column of eqnarray equal to twice \arraycolsep. I would prefer to make nocolsep the default, but it seems best that the normal behavior agrees with the usual documented behavior of L^{A}TEX.

This typeset documentation was generated by running L^{A}TEX on dotseqn.dtx, and dotseqn.sty is generated the same way. This is my experiment with singlefile dtx distribution.